OpenStudy (anonymous):

Simplify the expression completely. (-3d^4e^8)^3 (2d^2e^4)^3

OpenStudy (welshfella):

first simplify the 2 'halves' (-3d^4e^8)^3 = (-3d^4)^3*(e^8)^3 now use the index rule ( na^b)^c = n^c a^(bc)

OpenStudy (anonymous):

thx

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

so the first part comes to (-3)^3 d^(4*3) e^(8*3) = -27 d^12 e^24 now work out the decond part in a similar way then the 2 parts are multiplied

OpenStudy (welshfella):

recall the rule for multiplying indices a^n * a^m = a^(n + m)